# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

### v2

@@ Abstract
We elucidate the generic bifurcation behavior of local and global order in the short-range interacting nonreciprocal Ising model evolving under Glauber dynamics. We show that a critical magnitude of nearest-neighbor correlations within the respective lattices controls the emergence of coherent oscillations of global order as a result of frustration. Local order is maintained during these oscillations, implying non-trivial spatio-temporal correlations. Long-lived states emerge in the strong-interaction regime. The residence time in either of these states eventually diverges, giving rise to ordered non-equilibrium steady states via a saddle-node-infinite-period bifurcation. Our work provides a comprehensive microscopic understanding of collective phenomena in nonreciprocal matter without conservation laws.
The last decade saw a surge of interest in multicomponent systems with nonreciprocal interactions. Nonreciprocity is commonly found in active and non-equilibrium systems [1, 2, 3, 4, 5], such as predator-prey [6, 7] and many-body lattice [8, 9, 10, 11, 12, 13] systems, spin glasses [14, 15, 16], active solids with odd elasticity [17], cell-cell interactions [18], optically trapped nanoparticles [19, 20], complex plasma dynamics [21, 22, 23], mixtures of passive and active particles [24] as well as chemically active interacting particles [25, 26, 27, 28, 29], primitive metabolic cycles [30], and non-Hermitian quantum systems [31, 32, 33]. The importance of nonreciprocal interactions is also widely recognized in the social sciences [34, 35], neuroscience [36, 37], and robotics [38]. However, the underlying mechanisms and implications of nonreciprocity can vary substantially among these disciplines.
At the microscopic level, nonreciprocal interactions are closely linked to the violation of Newton’s third law and broken detailed-balance [21, 2]. On the collective level, nonreciprocally interacting systems can resist coarsening and instead self-organize into dynamic states with unique temporal and spatial patterns, including traveling and oscillatory states [39, 40, 41, 42, 43, 44, 45, 46]. Phenomenogically, such systems are typically described by a nonvariational coupling of Cahn-Hilliard models or Allen-Cahn models (i.e., the respective models B and A in [47]).
An important aspect of nonreciprocal systems is the presence of conservation laws. In the case of two nonvariationally coupled fields one may have no [48, 8, 16, 49], one [50, 51, 52], or two [2, 24, 40, 41, 46, 43, 53] conservation laws. Nonreciprocal systems without conservation laws are particularly relevant in the realm of chemical reaction networks, where it was shown that kinetic asymmetries between enzymes may lead to nonreciprocal interactions, resulting in spatio-temporal oscillations [28].
There is growing interest in a bottom-up understanding of non-conserved nonreciprocal systems, so far limited to phenomenological and mean-field theory [48, 8, 16]. These works revealed intriguing phenomena, such as large-scale Hopf instabilities [48, 8, 16], saddle-node bifurcations [49, 48], and hidden collective oscillations [16].
@@
Here, we go beyond mean-field reasoning and develop a thermodynamically consistent description of nonreciprocal binary systems without conservation laws, starting from a pair of coupled Ising models, which serves as a minimal model of a many-body system with short-range interactions. We incorporate nearest-neighbor correlations and consider both, global and local order. We explain why a critical magnitude of nearest-neighbor correlations controls the symmetry-breaking transition in the global order, in turn bounding the onset of coherent oscillations. We elucidate how the bifurcation behavior depends on inter-particle interactions, and highlight stark differences in the spatio-temporal dynamics of all-to-all (mean-field) versus short-range-interacting systems; the square and Bethe lattices display equivalent behavior that is strikingly different from the all-to-all lattice.
@@
Fundamentals.—Consider a pair of lattices denoted by $\mu=a,b$, as shown in Fig. 1a, each having a coordination number $z$ and periodic boundary conditions. On each lattice there are $N$ spins that can assume two states $\sigma^{\mu}_{i}=\pm 1$, with $i\in\{1,...,N\}$ enumerating the spin’s location. Each spin interacts with its $z$ nearest-neighbors on the same lattice and the spin at the equivalent position on the opposing lattice. The local interaction  [^1] The local interaction is not equal to the total energy of the system, since the interaction energy for the two lattices $a$ and $b$ is different. energy of spin $i$ on lattice $\mu$ can be written as (see also [48])
$E^{\mu}_{i}=-J_{\mu}\sigma^{\mu}_{i}\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}-K_{\mu}\sigma^{a}_{i}\sigma^{b}_{i},\$ (1)
@@
where $\langle i|j\rangle$ denotes a sum over nearest-neighbors $j$ of spin $i$. Throughout, we will express energies in units of $k_{\rm B}T$, where $T$ is the temperature of the heat bath. The parameters $J_{\mu}$ and $K_{\mu}$ denote the coupling within lattice $\mu$ and of the spins in $\mu$ with those on the opposing lattice, respectively. When $K_{a}\neq K_{b}$, equivalent spins on the opposing lattices interact nonreciprocally.
Figure 1: (a) Schematic of two Ising lattices $a$ and $b$ with coordination number $z{=}4$ and cross-coupling $(K_{a},K_{b})$. For $K_{a}\neq K_{b}$ the cross-coupling is nonreciprocal. (b) Eigenvalues $\lambda_{\pm}$ of the linear stability matrix (see Eq. (8)) as a function of the steady state local order $q_{s}$ for $K_{a}=-K_{b}=0.1$. When $q_{s}\geq q_{\rm crit}\equiv 1/(z-1)$ (black vertical line) the real parts of $\lambda_{\pm}$ are non-negative, resulting in coherent oscillations. (c-f) Temporal evolution of the magnetization (c; see Eq. (2)), (d) local order (see Eqs. (3)-(4)), and (e,f) local correlations $\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t)$ for $K_{a}=-K_{b}=K=0.1$ and $J_{a}=J_{b}=0.4$. Black dashed lines in (c,d) are obtained with Monte-Carlo simulations on the Bethe lattice (see [55] for details). In all panels we consider $z=4$.
@@
We consider single spin-flip Glauber dynamics [56]. Let $P(\boldsymbol{\sigma};t)$ be the probability at time $t$ to find the system in state $\boldsymbol{\sigma}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ which is governed by the master equation $dP(\boldsymbol{\sigma};t)/dt=\sum_{\mu}\sum_{i}[w^{\mu}_{i}(-\sigma^{\mu}_{i})P(\boldsymbol{\sigma}^{\prime}_{\mu,i};t){-}w^{\mu}_{i}(\sigma^{\mu}_{i})P(\boldsymbol{\sigma};t)]$, where $\boldsymbol{\sigma}^{\prime}_{\mu,i}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,-\sigma^{\mu}_{i},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ is a state which differs from state $\boldsymbol{\sigma}$ by one spin flip. The transition rates $w^{\mu}_{i}(\sigma^{\mu}_{i})$ to flip a spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$ are uniquely specified by limiting the interactions to nearest neighbors, imposing isotropy in position space, and requiring that for $K_{a}=K_{b}$ the transition rates obey detailed-balance. These physical restrictions then lead to the general result $w^{\mu}_{i}(\sigma^{\mu}_{i}){=}\left[1{-}\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\right]/2\tau$ [8, 48, 16], where $\Delta E^{\mu}_{i}=-2E^{\mu}_{i}$ is the change in energy on the $\mu=a\lor b$ lattice after spin conversion $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau$ is an intrinsic timescale to attempt a single spin-flip.
@@
We are interested in the temporal dynamics of global and local order parameters averaged over all spins. The magnetization or global order [57] is given by
$m^{\mu}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{\mu}_{i}(t)\rangle,$ (2)
@@
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Note that $m^{\mu}(t)\in[-1,1]$. The three local order parameters [57] are
$q^{\mu\mu}(t)$ $\equiv\frac{2}{zN}\sum_{i}\sum_{\langle i|j\rangle}\langle\sigma_{i}^{\mu}(t)\sigma_{j}^{\mu}(t)\rangle,$ (3)
$q^{ab}(t)$ $\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma_{i}^{a}(t)\sigma_{i}^{b}(t)\rangle,$ (4)
@@
with correlations $\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t)$. We distinguish between the local order within and between the lattices. The alignment of spin pairs within lattice $\mu$ is quantified by $q^{\mu\mu}(t)$, and $q^{ab}(t)$ measures the alignment of equivalent spins on opposing lattices. The normalization in Eq. (3) arises because $zN/2$ is the number of nearest neighbor pairs in a periodic lattice with coordination number $z$, which renders $q^{\mu\mu}(t),q^{ab}(t)\in[-1,1]$.
@@
Evolution equations.—Our first main result is an exact set of coupled differential equations for the order parameters in the thermodynamic limit $N\rightarrow\infty$, which reads (see [55] for a detailed derivation)
$\tau\frac{dm^{\mu}(t)}{dt}{+}m^{\mu}(t)$ ${=}\sum\nolimits_{l,n}\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$
$\tau\frac{dq^{\mu\mu}(t)}{dt}{+}2q^{\mu\mu}(t)$ ${=}\frac{2}{z}\!\sum\nolimits_{l,n}\!(2l{-}z)\mathcal{P}^{\mu}_{l,n}(t)\!\tanh(U^{\mu}_{l,n}),$ (5)
$\tau\frac{dq^{ab}(t)}{dt}{+}2q^{ab}(t)$ ${=}\!\sum\nolimits_{\mu}\sum\nolimits_{l,n}(2n{-}1)\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$
@@
where $\sum_{l,n}\equiv\sum_{l=0}^{z}\sum_{n=0}^{1}$ is a sum over all possible nearest neighbor spin values on the same ($l=0,\ldots,z$) and opposing ($n=0,1$) lattice, $U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$, and $\mathcal{P}^{\mu}_{l,n}(t)\in[0,1]$ is the time-dependent probability to select an up or down spin which has $l$ neighboring up spins on the same lattice, and $n$ neighboring up spins on the opposing lattice. The probability is normalized as $\sum_{l,n}\mathcal{P}^{\mu}_{l,n}(t){=}1$. Equations (5) are not yet closed; evaluating $\mathcal{P}^{\mu}_{l,n}(t)$ for an arbitrary lattice is a daunting combinatorial task, as it depends on microscopic details and therefore on an infinite hierarchy of order parameters.
@@
An accurate closed-form expression for $\mathcal{P}^{\mu}_{l,n}(t)$ can be obtained with the Bethe-Guggenheim (BG) approximation, where we assume perfect mixing of nearest neighbor spin pairs. Such an approximation is exact on loop-less lattices like the Bethe lattice [58], or large random graphs with fixed coordination number [59, 60]. As we show below, the spatio-temporal dynamics on these lattices captures the behavior on the square lattice. Note that a theory of the nonreciprocal Ising model on the fully-connected mean-field lattice has been constructed in [48], which we also report in Appendix A.
@@
We split $\mathcal{P}^{\mu}_{l,n}(t)$ in “up” and “down” spin contributions $\mathcal{P}^{\mu}_{l,n}(t)=\mathcal{P}^{\mu+}_{l,n}(t)+\mathcal{P}^{\mu-}_{l,n}(t)$, where, e.g., $\mathcal{P}^{\mu+}_{l,n}(t)\in[0,1]$ is the probability of flipping an up spin with $l$ up neighbors on the same and $n$ up neighbors on the opposing lattice, respectively. These probabilities read  [^2] See Supplementary Material at […] (omitting the explicit $t$-dependence on the r.h.s.)
@@
where $C^{z}_{l}{\equiv}\binom{z}{l}/2^{z+2}$, $\delta^{+}_{l}{=}l$, and $\delta^{-}_{l}{=}z-l$. The expression for $\mathcal{P}^{b\pm}_{l,n}(t)$ follows from Eq. (6) by replacing $m^{a}{\rightarrow}m^{b}$ and $q^{aa}{\rightarrow}q^{bb}$. Inserting Eq. (6) into Eqs. (5) yields a closed system of five coupled nonlinear differential equations.

### v3

# Local Order Controls the Onset of Oscillations in the Nonreciprocal Ising Model

@@ Abstract
We elucidate the generic bifurcation behavior of local and global order in the nonreciprocal Ising model evolving under Glauber dynamics. We show that a critical magnitude of nearest-neighbor correlations within the respective lattices controls the emergence of coherent oscillations of global order as a result of frustration. Local order is maintained during these oscillations, implying nontrivial spatiotemporal correlations. Long-lived states emerge in the strong-interaction regime. The residence time in either of these states eventually diverges, giving rise to ordered non-equilibrium trapped states and a loss of ergodic behavior via a saddle-node-infinite-period bifurcation. Our work provides a comprehensive microscopic understanding of the nonreciprocal Ising model beyond the mean-field approximation.
I Introduction
The last decade saw a surge of interest in many-body lattice systems with nonreciprocal interactions Guislain and Bertin (2024a, b); Seara et al. (2023); Collet et al. (2016); Collet (2014); Collet and Formentin (2019); Avni et al. (2023); Loos et al. (2023); Guislain and Bertin (2024c); Osat and Golestanian (2023); Osat et al. (2024). At the microscopic level, nonreciprocal interactions violate Newton’s third law and result in broken detailed balance Ivlev et al. (2015); Brauns and Marchetti (2024), thereby driving the system out of equilibrium. On the collective level, nonreciprocally interacting systems can resist coarsening and self-organize into dynamic states with unique spatiotemporal patterns, including traveling and oscillatory states te Vrugt et al. (2022); Frohoff-Hülsmann et al. (2021); Saha et al. (2020); John and Bär (2005); Frohoff-Hülsmann and Thiele (2021); Frohoff-Hülsmann et al. (2023); Mandal et al. (2024); Frohoff-Hülsmann et al. (2023); Greve et al. (2025). Phenomenologically, such systems are typically described using nonvariational couplings of Allen-Cahn models (for nonconserved dynamics) or Cahn-Hilliard models (for conserved dynamics), corresponding to models A and B, respectively, as outlined in Hohenberg and Halperin (1977).
By introducing two nonreciprocally coupled Ising lattices, various studies have shown under which conditions nonreciprocity induces temporal oscillations in the magnetization Avni et al. (2023); Guislain and Bertin (2024a); Seara et al. (2023); Guislain and Bertin (2024b); Collet et al. (2016); Collet (2014); Collet and Formentin (2019). These works revealed intriguing phenomena, such as Hopf instabilities Avni et al. (2023); Guislain and Bertin (2024a, b); Collet and Formentin (2019), saddle-node bifurcations Liu et al. (2023); Avni et al. (2023), and hidden collective oscillations Guislain and Bertin (2024b). However, so far these studies have been limited to phenomenological and mean-field theory, raising the question to what extent these results apply beyond their respective approximations.
@@ I Introduction
Here, we go beyond mean-field reasoning and explicitly incorporate nearest-neighbor correlations into a thermodynamically consistent description of two nonreciprocally coupled Ising models We consider both global and local order, and explain why a critical magnitude of nearest-neighbor correlations controls the symmetry-breaking transition in the global order, in turn bounding the onset of coherent oscillations. We elucidate how the bifurcation behavior depends on the interaction strength and highlight stark differences in the spatiotemporal dynamics of all-to-all (mean-field) versus short-range-interacting systems; the square and Bethe lattices display equivalent behavior that is strikingly different from the all-to-all lattice.
II Model
@@
Consider a pair of lattices denoted by $\mu=a,b$, as shown in Fig. 1a, each having a coordination number $z$ and periodic boundary conditions. On each lattice there are $N$ spins that can assume two states $\sigma^{\mu}_{i}=\pm 1$, with $i\in\{1,...,N\}$ enumerating the spin’s location. Each spin interacts with its $z$ nearest neighbors on the same lattice and the spin at the equivalent position on the opposing lattice. The local interaction  [^1] The local interaction is not equal to the total energy of the system, since the interaction energy for the two lattices $a$ and $b$ is different. energy of spin $i$ on lattice $\mu$ can be written as
$E^{\mu}_{i}=-J_{\mu}\sigma^{\mu}_{i}\sum_{\langle i|j\rangle}\sigma^{\mu}_{j}-K_{\mu}\sigma^{a}_{i}\sigma^{b}_{i},$ (1)
@@
where $\langle i|j\rangle$ denotes a sum over nearest-neighbors $j$ of spin $i$. Throughout, we express energies in units of $k_{\rm B}T$, where $T$ is the temperature of the heat bath. The parameter $J_{\mu}$ denotes the coupling within lattice $\mu$, and $K_{\mu}$ denotes the (directed) coupling between the spins in $\mu$ and those of the opposing lattice, respectively. When $K_{a}\neq K_{b}$, equivalent spins on the opposing lattices interact nonreciprocally.
We consider single spin-flip Glauber dynamics Glauber (1963). Let $P(\boldsymbol{\sigma};t)$ be the probability at time $t$ to find the system in state $\boldsymbol{\sigma}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ which is governed by the master equation
$\frac{{\rm d}P(\boldsymbol{\sigma};t)}{{\rm d}t}=\sum_{\mu,i}w^{\mu}_{i}(-\sigma^{\mu}_{i})P(\boldsymbol{\sigma}^{\prime}_{\mu,i};t){-}w^{\mu}_{i}(\sigma^{\mu}_{i})P(\boldsymbol{\sigma};t),$ (2)
@@
where $\boldsymbol{\sigma}^{\prime}_{\mu,i}=\{\sigma_{1}^{a},\sigma_{1}^{b},...,-\sigma^{\mu}_{i},...,\sigma_{N}^{a},\sigma_{N}^{b}\}$ is a state which differs from state $\boldsymbol{\sigma}$ by one spin flip. The transition rates $w^{\mu}_{i}(\sigma^{\mu}_{i})$ to flip a spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$ are uniquely specified by limiting the interactions to nearest neighbors, imposing isotropy in position space, and requiring that for $K_{a}=K_{b}$ the transition rates obey detailed balance. These physical restrictions then lead to the general result Guislain and Bertin (2024a, b)
$w^{\mu}_{i}(\sigma^{\mu}_{i})=\left[1-\tanh{\left(\Delta E^{\mu}_{i}/2\right)}\right]/2\tau,$ (3)
where $\Delta E^{\mu}_{i}=-2E^{\mu}_{i}$ is the change in energy on the $\mu=a,b$ lattice after spin conversion $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau$ is an intrinsic timescale to attempt a single spin-flip.
II.1 Global and local order parameters
@@
We are interested in the temporal dynamics of global and local order parameters averaged over all spins. The magnetization or global order Saito and Kubo (1976) is given by
$m^{\mu}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{\mu}_{i}(t)\rangle\ \in[-1,1],$ (4)
@@
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. The three local order parameters Saito and Kubo (1976) are
$q^{\mu\mu}(t)$ $\equiv\frac{2}{zN}\sum_{i=1}^{N}\sum_{\langle i|j\rangle}\langle\sigma_{i}^{\mu}(t)\sigma_{j}^{\mu}(t)\rangle\ \in[-1,1],$ (5)
$q^{ab}(t)$ $\equiv\frac{1}{N}\sum_{i=1}^{N}\langle\sigma_{i}^{a}(t)\sigma_{i}^{b}(t)\rangle\ \ \ \ \ \ \ \ \in[-1,1],$ (6)
with correlations
$\mathcal{C}^{\mu\nu}(t)\equiv q^{\mu\nu}(t)-m^{\mu}(t)m^{\nu}(t).$ (7)
@@
We distinguish between the local order within and between the lattices. The alignment of spin pairs within lattice $\mu$ is quantified by $q^{\mu\mu}(t)$, and $q^{ab}(t)$ (also known as the overlap Guislain and Bertin (2024d, 2023)) measures the alignment of equivalent spins between both lattices. The normalization in Eq. (5) arises because $zN/2$ is the number of nearest neighbor pairs in a periodic lattice with coordination number $z$.
II.2 Evolution equations beyond the mean-field approximation
@@
Our first main result is an exact set of coupled differential equations for the order parameters in the thermodynamic limit $N\rightarrow\infty$, which reads (see Appendix A for a detailed derivation)
$\!\!\!\!\tau\frac{{\rm d}m^{\mu}(t)}{{\rm d}t}{+}m^{\mu}(t)$ ${=}\sum_{l,n}\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$ (8)
$\!\!\!\!\tau\frac{{\rm d}q^{\mu\mu}(t)}{{\rm d}t}{+}2q^{\mu\mu}(t)$ ${=}\frac{2}{z}\!\sum_{l,n}(2l{-}z)\mathcal{P}^{\mu}_{l,n}(t)\tanh(U^{\mu}_{l,n}),$ (9)
$\!\!\!\!\tau\frac{{\rm d}q^{ab}(t)}{{\rm d}t}{+}2q^{ab}(t)$ ${=}\!\sum_{\mu}\!\sum_{l,n}(2n{-}1)\mathcal{P}^{\mu}_{l,n}(t)\!\tanh(U^{\mu}_{l,n}),$ (10)
where $\sum_{l,n}\equiv\sum_{l=0}^{z}\sum_{n=0}^{1}$ is a sum over all possible values of neighboring up spins on the same ($l\in\{0,\ldots,z\}$) and opposing ($n\in\{0,1\}$) lattice, and
$U^{\mu}_{l,n}\equiv[2l-z]J_{\mu}+[2n-1]K_{\mu}$ (11)
parameterizes the change in energy upon flipping a spin with such a local environment. Finally, $\mathcal{P}^{\mu}_{l,n}(t)\in[0,1]$ is the time-dependent probability of selecting an up or down spin that has $l$ neighboring up spins on the same lattice and $n$ neighboring up spins on the opposing lattice. The probability is normalized as
$\sum_{l,n}\mathcal{P}^{\mu}_{l,n}(t)=1.$ (12)
@@
Equations (8)-(10) are not yet closed; evaluating $\mathcal{P}^{\mu}_{l,n}(t)$ for an arbitrary lattice is a daunting combinatorial task, as it depends on microscopic details and, therefore, on an infinite hierarchy of order parameters. However, we can approximate $\mathcal{P}^{\mu}_{l,n}(t)$ to different levels of accuracy, which we do in the next section.
Note that evolution equations for the nonreciprocal Ising model on the fully connected mean-field lattice have been constructed in Avni et al. (2023), and are reported in Appendix E for completeness.
Figure 1: (a) Schematic of two Ising lattices $a$ and $b$ with coordination number $z{=}4$ and cross-coupling $(K_{a},K_{b})$. For $K_{a}\neq K_{b}$ the cross-coupling is nonreciprocal. (b) Eigenvalues $\lambda_{\pm}$ of the linear stability matrix [Eq. (22)] as a function of the steady-state local order $q_{s}$ for $K_{a}=-K_{b}=0.1$. When $q_{s}\geq q_{\rm crit}$ [black vertical line; Eq. (20)] the real parts of $\lambda_{\pm}$ are non-negative, resulting in coherent oscillations. (c-f) Temporal evolution of the magnetization [Eq. (4)] (c), local order [Eqs. (5)-(6)] (d), and local correlations [Eq. (7)] (e,f) for $K_{a}=-K_{b}=K=0.1$ and $J_{a}=J_{b}=0.4$. Black dashed lines in (c,d) are obtained with Monte-Carlo simulations on the Bethe lattice (see Appendix G for details). In all panels we consider $z=4$.
II.3 Bethe-Guggenheim approximation
@@
An accurate closed-form expression for $\mathcal{P}^{\mu}_{l,n}(t)$ can be obtained with the Bethe-Guggenheim (BG) approximation (or pair approximation), where we assume perfect mixing of nearest neighbor spin pairs. This approximation is exact on loopless lattices such as the Bethe lattice Blom (2023), or large random graphs with fixed coordination number Johnston and Plechác (1998); Dhar et al. (1997). As we show in Sec. V, the spatiotemporal dynamics on these lattices agrees, in contrast to that on a mean-field lattice, with the behavior on the square lattice. We split $\mathcal{P}^{\mu}_{l,n}(t)$ in “up” and “down” spin contributions
$\mathcal{P}^{\mu}_{l,n}(t)=\mathcal{P}^{\mu+}_{l,n}(t)+\mathcal{P}^{\mu-}_{l,n}(t),$ (13)
@@
where, e.g., $\mathcal{P}^{\mu+}_{l,n}(t)\in[0,1]$ is the probability of flipping an up spin with $l$ up neighbors on the same and $n$ up neighbors on the opposing lattice, respectively. These probabilities are derived in Appendix B and read (omitting the explicit $t$-dependence on the right-hand side)
@@
where $\delta^{+}_{l}=l$, $\delta^{-}_{l}=z-l$, and
$C^{z}_{l}\equiv\frac{1}{2^{z+2}}\binom{z}{l}.$ (15)
@@
The expression for $\mathcal{P}^{b\pm}_{l,n}(t)$ follows from Eq. (14) by interchanging $m^{a}\leftrightarrow m^{b}$ and $q^{aa}\leftrightarrow q^{bb}$. Inserting Eq. (14) into Eqs. (8)-(10) yields a closed system of five coupled nonlinear differential equations.
III Linear analysis and the Hopf bifurcation